### v1

Abstract
Recently, Large Language Models (LLMs) have dominated much of the artificial intelligence scene with their ability to process and generate natural languages. However, the majority of LLM research and development remains English-centric, leaving low-resource languages such as those in the Southeast Asian (SEA) region under-represented. To address this representation gap, we introduce Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT, two cutting-edge multilingual LLMs designed for SEA languages. The SEA-LION family of LLMs supports 11 SEA languages, namely English, Chinese, Indonesian, Vietnamese, Malay, Thai, Burmese, Lao, Filipino, Tamil, and Khmer. Our work leverages large-scale multilingual continued pre-training with a comprehensive post-training regime involving multiple stages of instruction fine-tuning, alignment, and model merging. Evaluation results on multilingual benchmarks indicate that our models achieve state-of-the-art performance across LLMs supporting SEA languages. We open-source the models [^1] SEA-LION Models Collection to benefit the wider SEA community.
SEA-LION: Southeast Asian Languages in One Network
1 Introduction
Large language models (LLMs) have significantly transformed the field of natural language processing, achieving remarkable performance in text generation, summarization and sentiment analysis (Brown et al., 2020; OpenAI, 2023; Dubey et al., 2024; Rivière et al., 2024; Zhang et al., 2024b; Yeo et al., 2024). Despite the impressive capabilities of LLMs, the vast majority of them still are very much English-centric (Wendler et al., 2024; Zhong et al., 2024).
Unfortunately, this situation has led LLMs in regions with many under-represented languages such as Southeast Asia (SEA) to suffer. Languages with lower resources, such as Filipino, Lao, Burmese and Khmer in the SEA region, are not supported by many open-source English-centric LLMs such as Llama (Dubey et al., 2024) and Olmo (Groeneveld et al., 2024). This raises a pressing need to mitigate this language resource and representation gap between English and SEA languages.
Recently, there have been many attempts to create multilingual LLMs in an open-source manner. For instance, BLOOM (Scao et al., 2022) was a project aimed at increasing multilingual presence in open-source LLMs by supporting 46 natural languages. Popular LLM families such as Llama (Dubey et al., 2024), Gemma (Rivière et al., 2024) and Qwen (Yang et al., 2024a) have also introduced multilingual LLMs for their latest iteration. During our evaluations, we found that the performance of these models is acceptable in the general case, i.e., if we consider evaluation benchmarks formulated from English datasets, but we observe that the performance degrades on SEA-specific benchmarks.
Moreover, researchers have also introduced LLMs such as SeaLLMs Nguyen et al. (2024); Zhang et al. (2024a) and Sailor Dou et al. (2024) to specifically address the LLM gap in SEA languages. However, the performance of these models is less than ideal for languages such as Thai or Tamil [^2] Tamil is one of the official languages in Singapore. It is also spoken in other areas in the SEA region, such as Malaysia. (10X et al., 2024; AI Products Team, 2024).
In this paper, we address the issues by proposing a robust open-source Southeast Asian model with data transparency for reproducibility, namely SEA-LION – a family of LLMs CPT and fine-tuned on Llama-3.1-8B-Instruct and Gemma-2-9B with a focus on SEA languages. To tackle the performance problem, we utilize 200 billion English, code and SEA languages tokens as well as 16.8 million English and SEA languages instruction and answer pairs for CPT and post-training steps respectively, to achieve a significant improvement in SEA languages. In order to allow our models to be used by everyone without restrictions, we release our models under a fully open MIT license. We benchmark our models against the SEA-HELM [^3] SEA-HELM Leaderboard (Susanto et al., 2025) and Open LLM Leaderboard [^4] Open LLM Leaderboard with other LLMs of similar sizes in Southeast Asia like Sailor 2 (Team, 2024) and SeaLLMs 3 (Zhang et al., 2024a) where our models achieve state-of-the-art performances. We summarize the contribution of our paper as follows.
    We released two LLMs, Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT, that are meticulously trained to accurately represent the unique linguistic diversity of SEA languages.
    We also provide in-depth insights in this paper into our end-to-end training workflow to benefit the community developing multilingual LLMs.
2 Continued pre-training (CPT)
2.1 Pre-training data
The CPT data consists of a curated set of English, multilingual and code corpora from several open source repositories like Dolma (Soldaini et al., 2024), FineWeb (Penedo et al., 2024), the-stackv2 (Lozhkov et al., 2024), SEA-LION-Pile (AI Singapore, 2023), SEA-LION-Pile-v2 (AI Singapore, 2025), as well as documents from CommonCrawl (CommonCrawl, 2024) and from the public domain such as Wikipedia (Foundation, 2024). For SEA-LION-Pilev2, we filter CommonCrawl WARC data for documents in SEA languages (i.e., Burmese, Simplified Chinese, Indonesian, Khmer, Lao, Malay, Filipino, Tamil, Thai and Vietnamese) using the pretrained fasttext language classifier (Joulin et al., 2017). A document is retained if the language code reported in its metadata matches with that of one of the aforementioned SEA languages. Additionally, we further clean up the data with Trafilatura (Barbaresi, 2021).
To determine the best dataset ratio between SEA languages, code and English for the CPT process, we perform a series of small-scale CPT experiments each with a training budget of 10B tokens and varying proportions of English, code and SEA language data. We settled on an optimal data mix ratio of 55% SEA languages, 25% English and 20% code tokens for a budget of 200B tokens. For a detailed breakdown of the token count by languages, please refer to the model card. [^5] Gemma-SEA-LION-v3-9B Model Card
2.2 CPT process
Model selection. For the models to CPT from, we choose Llama-3.1-8B-Instruct (Dubey et al., 2024) and Gemma-2-9B (Rivière et al., 2024).
Training setup. Following previous works Dou et al. (2024), we use BPE-Dropout (Provilkov et al., 2020) to increase the performance and robustness of the training. We use a Warmup-Stable-Decay (WSD) (Hu et al., 2024) scheduler with warm-up and cooldown phases each representing 10% of the entire training budget. We use the AdamW (Loshchilov and Hutter, 2019) optimiser with the maximum learning rate (LR) set to ${1e^{-5}}$ and the final LR after cooldown is ${1e^{-7}}$. Following Wortsman et al. (2024), we set epsilon to ${1e^{-15}}$. We use Composer (Team, 2021) and LLM Foundry (Team, 2022) for distributed training using Fully Sharded Data Parallel (Zhao et al., 2023) on a cluster of eight nodes of the p5.48xlarge instance from Amazon Web Services (AWS). The total training duration was approximately 6 days and 10 days for the Llama 3.1 and Gemma 2 models, respectively. In this paper, we refer to the post-CPT models as Llama-SEA-LION-v3-8B and Gemma-SEA-LION-v3-9B for the Llama 3.1 and Gemma 2 continued pre-trained models respectively.
3 Post-training
3.1 Post-training data
The post-training data for instruction fine-tuning consists of Infinity-Instruct [Foundation and Chat] (Beijing Academy of Artificial Intelligence, 2024), OpenMath-Instruct 2 (Toshniwal et al., 2024) and our own SEA-Instruct. In particular, SEA-Instruct consists of multiple open-source instruction datasets, a synthetically generated dataset following the Magpie (Xu et al., 2024) template, and hand-crafted datasets collected from native Southeast Asians. The full detail of the SEA-Instruct and SEA-Preference dataset can be found in the model card. [^6] Gemma-SEA-LION-v3-9B-IT Model Card
Figure 1: Training process of Llama-SEA-LION-v3-8B-IT (Section 3.2.1). The post-training process consists of 2 stages of instruction fine-tuning, an alignment stage and multiple merge stages. Dotted lines denote a merge stage and solid lines denote an alignment stage.
3.2 Post-training process
We use LLaMaFactory (Zheng et al., 2024b) with DeepSpeed (Rasley et al., 2020) for all Instruction Fine Tuning (IFT) and alignment steps. All IFT stages are performed using full model fine-tuning where the models are from the previous step (Section 2.2) and existing models. We use MergeKit (Goddard et al., 2024) with a value of 1 for weight and density parameters for all merge steps. Models selected for merging are selected empirically, based on the openness of model licenses, the suitability for merging and performance.
3.2.1 Llama-SEA-LION-v3-8B-IT
Stage 1 IFT As shown in Figure 1, we started off the post-training phase with IFT of Llama-SEA-LION-v3-8B with the Infinity Instruct (Foundation) (Beijing Academy of Artificial Intelligence, 2024) and OpenMathInstruct2 Toshniwal et al. (2024) datasets. Both datasets contain approximately 9.5 million instruction pairs, primarily in English and centered around reasoning, math, and code. We refer to the model at this stage as Stage-1-Llama.
Stage 2 IFT We performed a second round of IFT using the SEA-Instruct dataset, which consists of approximately 7.3 million instruction pairs, of which 5 million instruction pairs are generated using the Gemma-2-27B-Instruct (Rivière et al., 2024) model and the Qwen2.5-32B-Instruct model Yang et al. (2024a) in SEA languages. The remaining are English language instruction pairs from the Infinity-Instruct (Chat) Beijing Academy of Artificial Intelligence (2024) dataset. We refer to the model at this stage as Stage-2-Llama.
First merge After finishing the IFT stages, we performed the first of a series of merges by merging Stage-1-Llama and Stage-2-Llama into the Llama-SEA-LION-v3-8B using the DARE TIES (Yu et al., 2024; Ilharco et al., 2023) method. We refer to the model at this stage as Merge-1-Llama.
Second merge In order to mitigate catastrophic forgetting due to the fine-tuning process (Alexandrov et al., 2024), we performed the second round of merge by merging top-performing instruction-tuned models that share the Llama 3.1 lineage. We merge the original Llama-3.1-8B-Instruct, Llama3-8B-SEA-LION-v2.1-Instruct (SEA-LION Team, 2024), and SuperNova-Lite (Arcee-AI, 2024) into Merge-1-Llama using the Consensus TA (Wang et al., 2024b; Ilharco et al., 2023) merge method. We refer to the model at this stage as Merge-2-Llama.
Helpfulness and preference alignment We performed one round of alignment on Merge-2-Llama using SimPO (Meng et al., 2024) with the SEA-Preference dataset. We refer to the model at this stage as Aligned-SimPO-Llama.
Final merge Lastly, we perform a merge using the DELLA-Linear merge. With the original Llama-3.1-8B-Instruct model as the base for merging, we merge in Merge-2-Llama and Aligned-SimPO-Llama to produce the final model, Llama-SEA-LION-v3-9B-IT.
3.2.2 Gemma-SEA-LION-v3-9B-IT
Figure 2: Training process of Gemma-SEA-LION-v3-9B-IT (Section 3.2.2). The post-training process comprises two stages of instruction fine-tuning, an alignment stage, and multiple merge stages. Dotted lines denote a merge stage and solid lines denote an alignment stage.
Stage 1 and Stage 2 IFT Similar to the Llama-SEA-LION-v3-8B-IT, we started off the post-training phase with both stages of IFT using the same datasets on the Gemma-2-9B model (Rivière et al., 2024). We refer to both models at stage 1 and stage 2 as Stage-1-Gemma and Stage-2-Gemma, respectively.
First merge We merge the Gemma-2-9B-IT (Rivière et al., 2024) and Stage-2-Gemma into Gemma-2-9B using the DELLA Linear method. We refer to the model at this stage as the Merge-1-Gemma.
Helpfulness and preference alignment Using the Merge-1-Gemma as the base model, we performed one round of alignment using SimPO with the SEA-Preference dataset. We refer to the model at this stage as the Aligned-SimPO-Gemma.
Final merge Finally, using the Gemma-2-9B model as the base model, we merged Merge-1-Gemma, FuseChat Gemma-2-9B-Instruct (Yang et al., 2024b), Gemma-SEA-LION-v3-9B, and Aligned-SimPO-Gemma into it to produce the final model Gemma-SEA-LION-v3-9B-IT.
3.3 Discussion
This post-training workflow emphasizes the careful balance between general capabilities, SEA-specific linguistic fluency, and natural conversational abilities. Each step in the workflow is designed to progressively refine the model, ensuring it meets the diverse needs of users in the Southeast Asian region.
The entire post-training process for Gemma-SEA-LION-v3-9B-IT and Llama-SEA-LION-v3-8B-IT took approximately 1350 and 1024 GPU hours respectively on eight H100 GPUs. To make the training efficient, all post-training steps utilize Liger Kernel (Hsu et al., 2024) for substantial memory savings of approximately 60%.
4 Experimental Setup and Results
SEA-HELM
NLU, NLG, NLR, NLI Instruction Following MTBench
Models Average ID VI TH TA ID VI TH ID VI TH
SeaLLMs-v3-7B-Chat 39.19 42.72 48.50 42.59 12.06 57.14 53.33 47.00 59.81 65.24 56.59
Llama-3.1-8B-Instruct 41.48 51.50 51.31 45.32 15.40 77.14 75.24 63.00 56.38 57.59 54.34
Sailor2-8B-Chat 43.13 48.98 48.01 45.44 28.29 49.52 45.71 40.00 69.76 66.97 73.94
Qwen2.5-7B-Instruct 44.58 60.28 53.46 53.43 21.03 81.90 69.52 66.00 65.66 66.80 68.71
Gemma-2-9B-IT 55.33 64.04 59.86 57.22 52.28 88.57 78.10 71.00 68.78 68.37 73.51
Stage-1-Llama 50.76 51.84 51.83 46.23 27.53 69.52 73.33 59.00 42.74 46.41 46.46
Stage-2-Llama 59.49 53.87 55.18 50.92 44.80 77.14 76.19 67.00 50.90 53.72 46.97
Merge-1-Llama 59.36 56.73 56.82 51.71 46.63 81.90 82.86 67.00 57.04 54.01 50.28
Merge-2-Llama 58.01 59.19 52.63 51.89 35.40 87.62 80.95 78.00 56.38 59.32 58.86
Aligned-SimPO-Llama 51.30 54.86 51.69 46.77 26.40 82.86 80.00 68.00 68.20 64.68 64.92
Llama-SEA-LION-v3-8B-IT 61.84 60.50 61.48 55.92 43.61 84.76 85.71 76.00 62.65 68.32 65.13
Stage-1-Gemma 56.56 55.06 54.51 51.96 42.74 66.67 74.29 61.00 47.35 47.26 55.05
Stage-2-Gemma 66.66 64.10 61.76 56.90 57.85 89.52 82.86 76.00 60.54 58.93 58.76
Merge-1-Gemma 69.26 66.25 64.95 59.74 60.41 89.52 91.43 82.00 66.45 64.47 65.00
Aligned-SimPO-Gemma 69.37 65.69 65.47 59.51 57.38 86.67 88.57 78.00 68.89 73.67 73.51
Gemma-SEA-LION-v3-9B-IT 69.35 66.26 64.93 59.23 58.82 94.29 88.57 78.00 65.85 73.27 69.07
Table 1: SEA-HELM multilingual benchmark on NLU, NLG, NLR, NLI, instruction following and multi-turn chat on instruct models of similar sizes.
Open LLM Leaderboard
Models Average MMLU-PRO BBH GPQA MATH Lvl 5 IFEval (EN) MUSR
Sailor2-8B-Chat 16.37 27.93 27.15 3.47 0.00 37.49 2.19
SeaLLMs-v3-7B-Chat 22.49 33.93 24.37 7.27 15.86 44.10 9.38
Llama-3.1-8B-Instruct 27.88 29.36 26.10 10.63 17.45 77.03 6.75
Qwen2.5-7B-Instruct 27.93 37.00 34.72 10.18 0.00 76.34 9.34
Gemma-2-9B-IT 28.86 31.95 42.14 14.77 0.23 74.36 9.74
Stage-1-Llama 24.51 25.87 26.32 7.83 19.26 62.89 4.88
Stage-2-Llama 27.75 28.10 24.64 7.72 19.56 78.78 7.74
Merge-1-Llama 27.49 27.47 26.22 8.28 19.79 76.16 7.04
Merge-2-Llama 29.96 29.92 28.78 9.96 19.94 82.61 8.54
Aligned-SimPO-Llama 30.58 30.84 34.31 8.39 26.59 75.76 7.61
Llama-SEA-LION-v3-8B-IT 30.39 31.01 29.47 10.40 22.58 80.35 8.54
Stage-1-Gemma 29.88 33.34 38.51 10.74 24.17 56.87 15.66
Stage-2-Gemma 33.48 34.67 36.06 11.74 20.77 83.00 14.61
Merge-1-Gemma 35.15 36.22 41.42 15.32 26.28 82.09 9.59
Aligned-SimPO-Gemma 35.31 37.65 42.38 14.99 27.79 80.23 8.82
Gemma-SEA-LION-v3-9B-IT 35.43 36.94 43.39 15.10 24.24 81.85 11.07
Table 2: Open LLM Leaderboard benchmarks across different instruct models of similar sizes.
4.1 Setup
For the evaluation, we compared our models against well-known LLMs, such as SeaLLMsv3 Zhang et al. (2024a), Sailorv2 Team (2024), Qwen 2.5 Yang et al. (2024a), Gemma 2 Rivière et al. (2024) and Llama 3.1 Dubey et al. (2024), where the parameters of those models are less than 10 billion parameters, similar to our models. The evaluations are split into two areas as follows.
Multilingual performance. We evaluated the multilingual performance of each LLM using the SEA-HELM Leaderboard (Leong et al., 2023; Susanto et al., 2025). Due to the lack of proper benchmarks for low-resource languages (e.g. Lao, Khmer, Filipino), we only benchmarked the languages covered in the SEA-HELM leaderboard, which are Indonesian, Tamil, Thai, and Vietnames [^7] SEA-HELM has been updated since this paper was written to include Filipino. We selected SEA-HELM because the design choice of this benchmark reflects the performance of SEA culture and knowledge the most compared with other existing benchmarks DAMO-NLP-SG (2024); Lovenia et al. (2024); Wang et al. (2024a). We used the evaluation code from the official website [^8] SEA-HELM Repository without any changes.
English performance. We evaluated the English performance of the models using the Open LLM Leaderboard (HuggingFace, 2024). The leaderboard consists of six benchmarks, IFEval Zhou et al. (2023), Big Bench Hard (Suzgun et al., 2023), MATH (Hendrycks et al., 2021), GPQA (Rein et al., 2023), MuSR (Sprague et al., 2024) and MMLU-PRO (Wang et al., 2024c).
4.2 Results
Multilingual performance. As shown in Table 1, the SEA-HELM benchmark performance demonstrates that our instruct models, Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT, attain competitive performance in SEA languages, with Gemma-SEA-LION-v3-9B-IT achieving one of the highest average performances. Both Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT outperform other SEA languages-focused LLMs, such as Sailor2-8B-Chat and SEALLMs-v3-7B-Chat, with an average score of 69.35 across all the languages covered by the SEA-HELM benchmark apart from the SEA-MTBench tasks.
English performance. The Open LLM Leaderboard performance is shown in Table 2. Both Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT performed competitively in English language, math, and reasoning tasks, with Gemma-SEA-LION-v3-9B-IT achieving the highest average score of 35.43.
4.3 Performance Analysis
Continued Pre-Training The CPT stage is primarily focused on gaining SEA languages capabilities and knowledge. For the purpose of comparison against base and CPT models, we observed a 6.05 and 7.19 average SEA-HELM performance increase over the Meta-Llama-3.1-8B and Gemma-2-9B for Llama-SEA-LION-v3-8B and Gemma-SEA-LION-v3-9B, respectively. We observed a much larger average increase with instruction following capabilities in particular, which we attribute to the fact that our CPT models are trained from the instruct models rather than from the base models. Both CPT models also managed to perform competitively against the Meta-Llama-3.1-8B and Gemma-2-9B base models on the Open LLM Leaderboard benchmarks. This indicates that our choice of retraining with a proportion of 25% English tokens has been beneficial in mitigating catastrophic forgetting which has been shown to stem from CPT (Zheng et al., 2024a).
As shown in Table 1, we chose Gemma since it is the most performant on multilingual benchmarks. However, we also show that our framework generalizes for every LLM by applying our framework on Llama3.1; although the performance of Llama3.1 is lower than Qwen or Sailor, we can still improve it to outperform all of them. Note that we have shown the full performance score of our CPT models and other base models in Appendix A.1.
Stage 1: English instruction fine tuning In Stage 1 IFT, the focus is predominantly on gaining general capabilities in math, code and general instruction following in the English language. Although our CPT models are based off of the instruct versions of Llama-3.1-8B, the CPT process has eroded the instruction following capabilities (See Table 2). We observe an increase of 3.86 and 9.72 for Stage-1-Llama and Stage-1-Gemma respectively in English instruction following capabilities on the IFEval benchmark. We also observe an average increase of 7.9 for Stage-1-Llama and 7.47 for Stage-1-Gemma for the SEA-HELM benchmark.
Stage 2: Multilingual instruction fine tuning In Stage 2 IFT, the focus is on multilingual and reasoning capabilities. By instruction fine tuning on SEA languages and higher complexity English instruction pairs, the Stage 2 models saw an average increase of 8.73 for Stage-2-Llama and 10.1 for Stage-2-Gemma over Stage 1 models on the SEA-HELM benchmark.
Merge 1: Combining Stage 1 and Stage 2 Despite the significant gains observed in Stage 1 and 2, we observed that the effects of catastrophic forgetting from earlier stages could still be observed after Stage 2. In order to mitigate this, we merge Stage 1 and Stage 2 models into the CPT model, after which we we observed an average increase of 2.6 for Merge-1-Gemma. We also observed an increase across all SEA-HELM benchmark tasks for Merge-1-Llama.
Merge 2: Incorporating instruct models To reintroduce helpfulness, relevance and informativeness of responses observed in Llama 3.1 and Gemma 2 models, we perform further merges of open-source instruct models. While we observed significant increases in MT-Bench benchmark scores for Vietnamese and Thai, we also observed a slight degradation of average SEA-HELM performance as well as a slight degradation of Indonesian MTBench scores, which we view as acceptable tradeoffs for the significant performance increases in Vietnamese and Thai. Note that since Merge-1-Gemma already demonstrates great performance across the SEA-HELM benchmark, we choose to skip this step for the Gemma model.
Alignment steps In the alignment step to align the models to human preference, we prioritize the SEA MTBench performance over the other SEA-HELM benchmark tasks. We observed a broad increase in SEA MTBench performances across all languages for both models. However, this comes with minor degradation of instruction following capabilities and overall Indonesian SEA-HELM performance.
The alignment step significantly pushes the model towards longer, more helpful and sensitive responses, but also compromises performance in more task-specific benchmarks and instruction following in some languages, which we try to resolve in the next step.
Final merge: Combining aligned models To compensate for the capability degradation in the previous steps, we merge Merge-2-Llama and Merge-1-Gemma with Aligned-SimPO-Llama and Aligned-SimPO-Gemma and various open sourced pre-trained models describe in sections 3.2.1 and 3.2.2 for their respective model families. For Llama-SEA-LION-v3-8B-IT, we observed a significant increase in average SEA-HELM performance (61.84) from the alignment stage (51.30), mainly from the increase in performance for the core tasks in SEA-HELM. This performance increase demonstrates the value of empirical selection of pre-trained models to be merged in based on each model’s strengths and weaknesses to produce a far superior model. For Gemma-SEA-LION-v3-9B-IT, it easily achieves higher performance compared to the Llama-SEA-LION-v3-8B-IT with fewer post training steps. We attribute this performance to the high performance of the base Gemma 2 model and also to the larger vocabulary size which have been demonstrated (Takase et al., 2024) to produce better models.
5 Conclusion
Despite the sizable population and language diversity in Southeast Asia, there remains a scarcity of resources and accurate linguistic and cultural representation with open source LLMs. In this paper, we introduce Llama-SEA-LION-v3-8B-IT and Gemma-SEA-LION-v3-9B-IT, two multilingual LLMs comprehensively trained to achieve state-of-the-art performances in SEA languages, based on the Llama and Gemma family of LLMs. SEA-LION represents the next advancement in the development of LLMs that explicitly supports SEA languages. Both models are fully open-source and available for commercial use to increase accessibility and innovation in multilingual LLMs in Southeast Asia.
Appendix A Appendix
A.1 Benchmarks for Continued Pretrained Models
SEA-HELM
NLU, NLG, NLR, NLI Instruction Following
Models Average ID VI TH TA ID VI TH
Meta-Llama-3.1-8B 35.37 42.33 40.67 35.13 38.88 16.19 19.05 9.00
SeaLLMs-v3-7B 37.04 44.79 48.29 43.53 27.45 26.67 35.24 26.00
Gemma-2-9B 41.48 47.65 43.28 42.00 53.26 4.76 3.81 10.00
Qwen2.5-7B 41.98 51.63 52.17 46.55 36.60 31.43 36.19 30.00
Sailor2-8B 42.62 53.23 47.33 46.64 45.04 30.48 30.48 35.00
Llama-SEA-LION-v3-8B 41.42 44.98 46.25 42.79 43.03 25.71 32.38 23.00
Gemma-SEA-LION-v3-9B 48.67 57.16 49.39 47.16 60.56 25.71 20.00 27.00
Table 3: SEA-HELM multilingual benchmark on NLU, NLG, NLR, NLI and instruction following on base and continued pre-trained models of similar sizes.
Open LLM Leaderboard
Models Average MMLU-PRO BBH GPQA MATH Lvl 5 IFEval (EN) MUSR
Meta-Llama-3.1-8B 13.9 24.95 25.29 6.32 5.14 12.7 8.98
Sailor2-8B 17.71 25.74 27.62 4.87 7.02 21.95 19.03
Gemma-2-9B 21.15 34.48 34.1 10.51 13.14 20.4 14.3
SeaLLMs-v3-7B 24.00 35.71 34.57 9.28 18.81 32.94 12.68
Qwen2.5-7B 24.99 37.39 35.81 9.96 18.88 33.74 14.14
Llama-SEA-LION-v3-8B 16.61 27.6 26.04 7.49 9.89 16.56 12.07
Gemma-SEA-LION-v3-9B 22.41 32.78 37.24 10.29 9.89 30.12 14.11
Table 4: Open LLM Leaderboard benchmarks across different continued pre-trained models of similar sizes.